# SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models

### v1

@@ 4.3.2. One-Step SQL Generation
For scenarios where high accuracy in SQL generation is required and the usage domain is relatively stable, we apply a one-step approach that leverages supervised fine-tuning (SFT) of LLMs to allow direct generation of SQL statements from user inputs. Notably, we use Hunyuan-70B-SQL as the generation model, which is specifically designed for SQL generation by Tencent. While data quality is critical to the accuracy of the fine-tuned model, we develop an automated data pipeline which includes two critical steps: data preparation and data refinement.
Data Preparation. In the data preparation phase, we begin by selecting a batch of natural language requirements from business data, along with their ground-truth SQL generation results. Given that these requirements are highly standardized and contain complete contextual information, relying solely on this data poses challenges in ensuring the model’s performance in scenarios with ambiguous information.
To address this, we first employ a LLM to process each query, preserving the original intent while minimizing the information provided. This approach generates a set of semantically ambiguous inputs. In addition, we prepare relevant domain knowledge to serve as a reference for these inputs. We leverage knowledge retrieval techniques outlined in Section 4.1 to extract supplementary knowledge related to the original queries, while also enabling the LLM to supplement the rules it relies on during the information reduction process. An example of data preparation without ground-truth results is as below,
Finally, the dataset we use for fine-tuning comprises the reduced queries, the supplementary contextual knowledge, and the ground-truth SQL outputs. This comprehensive approach enhances the model’s ability to perform effectively in diverse and ambiguous scenarios.
Data Refinement. After fine-tuning on the prepared data, we further introduce an adaptive data refinement method to improve the accuracy of the generation model. We develop two strategies to improve the data quality based on error types and user feedbacks. The first strategy involves automatically assessing the model’s generated results and categorizing the failed examples into different types, such as time processing errors, unit conversion errors. Based on this information, we will prioritize queries that address these needs in the next data preparation phase. The second strategy involves conducting small-scale user testing, where users provide simple evaluations of the SQL queries generated by the model. Examples that receive low ratings will be annotated and added to the new training set. After evaluating the current model with these two strategies, we will iterate on the data preparation phase and fine-tune the model again. This process will continue until the model’s accuracy meets the requirements or until we reach the budget limit set by our clients.
@@ 4.3.3. Two-Step SQL Generation
To meet the demand for a method with strong generalization capabilities and low deployment costs across multiple SQL generation scenarios, we propose a novel two-step generation paradigm. This paradigm breaks down the SQL generation task into two stages: the Semantic Model stage and SQL Model stage.
In the Semantic Model stage, we first structure the user’s input. In practical NL2SQL tasks, user inputs are often difficult to control, and incomplete or ambiguous inputs can directly affect the SQL generation accuracy of ChatBI systems. This, in turn, reduces user experience and raises the user threshold. Therefore, we enhance the semantic information of the inputs before generating SQL from user inputs. For a user query , we first perform a retrieval from a knowledge base that includes explanations of domain-specific terms. Notably, our knowledge management module (Section 4.1) allows for plug-and-play management, which boosts efficiency of deployment in industrial applications. The retrieved domain knowledge, along with , is then processed by a LLM-based Semantic Model for semantic analysis and understanding, resulting in the generation of structured text. In our implementation, the structured text takes various forms, such as question rewriting, JSON, and API formats. Additionally, during the Semantic Model stage, we employ automated demonstration selection and Chain-of-Thought (CoT) prompting techniques. These strategies greatly enhance the accuracy and generalization capability of the Semantic Model for different types of questions.
In the SQL Model stage, we use another LLM as the SQL generation model, specifically Hunyuan-70B-SQL. The enhanced structured information from the Semantic Model serves as the input for the final SQL generation. As the structured input provides ample context for SQL generation, the SQL Model can effortlessly generate SQL dialects tailored to various database systems from the same input. By structuring and supplementing user questions before generating SQL, the two-step generation paradigm ensures the accuracy of SQL generation, and meanwhile facilitates easy migration across different industrial scenarios. This approach significantly reduces the training costs and improves practical deployment efficiency.
| Category | Example | Description |
| Descriptive Requests | Which day had the highest revenue? | This category is a general term for basic analysis problems, including correlation, extreme value, difference analysis, etc. |
| Predictive Requests | Forecast failure rate for the next week. | This category include future trend analysis or label prediction of unseen data. |
| Diagnostic Requests | Identify anomalies in the sales revenue. | This category includes anomaly detection and metric attribution for numerical data, as well as drill-down for categorical data. |
| Prescriptive Requests | How to deal with abnormal values of production volume? | This category helps interpret data insights and proposes actionable solutions based on the analysis. |
| Complex Requests | Forecast the net income for next week and calculate the expected year-over-year change. | This category involves executing multiple sub-tasks and there are dependencies between the execution of these sub-tasks. |
4.4. Data Insight
Data insights are essential to ChatBI systems, as they enable users to extract valuable information from raw data. This functionality allows users not only to understand the current state of affairs but also forecast future trends. Data insights go beyond traditional descriptive analytics; they reveal hidden patterns in the data and provide deep understanding, including root cause analysis of issues, predictive modeling for the future, and actionable recommendations based on the analysis. Such in-depth analysis is crucial for businesses because it helps decision-makers make informed choices, thus driving business growth. SiriusBI leverages the capabilities of LLMs to refine the functionalities of data insights, which greatly improves the user experience. The data insight module of SiriusBI is designed to address various problems as detailed in Table 1.
4.4.1. Main Workflow
The primary workflow for data insights in SiriusBI includes the following steps: automatic data preparation, complex task planning, invocation of insight tools, and generation of insight reports. A more detailed process flow will be introduced as follows.
Automatic Data Preparation. The insight tool executes three sequential steps during automatic data preparation: intent rewriting, Text2DSL, and metric-dimension recommendation. Initially, intent rewriting categorizes and refines the given insight question into a more specific description. Subsequently, Text2DSL converts the question into a Domain-Specific Language (DSL) corresponding to SQL queries. By inputting the schema information of the data table along with the insight question, this part outputs structured DSL language that includes essential columns for data insight. Finally, the schema of the data table, the DSL output, and the insight question are input into the metric-dimension recommendation part, which suggests additional relevant data columns that could fulfill the insight requirements based on the type of insight task.
Complex Task Planning. The Data Insight module is capable of planning complex tasks, in which a question may contains multiple interdependent requirements. It employs an internal planner that decompose these tasks into manageable sub-tasks and iteratively refines the plan based on the execution results of the previous sub-task. For instance, consider a scenario where the module is tasked with calculating the month-over-month growth of net new holdings for July 8, 2024, using a business dataset. Since the data for that specific date is not available in the data table, the execution process must first include a predictive analysis to estimate the value for July 8 before proceeding with the growth calculation.
Insight Tool Execution. The capabilities of the data insight module are supported by specialized insight tools. Depending on the specific tasks, the insight tools in SiriusBI primarily include:
Code Generation. For straightforward queries, such as retrieving a specific piece of data, SiriusBI utilizes a LLM to generate the necessary code, which is then executed in a secure sandbox environment.
Data Modeling. For specialized tasks such as prediction and anomaly detection, SiriusBI develops specialized machine learning models and utilizes automated tools for hyperparameter optimization, providing users with accurate output support.
Visualization. For visualization needs, SiriusBI uses a LLM to construct configuration files of visualization settings, which are then used by underlying visualization tools to automatically generate charts.
Knowledge Retrieval. For questions that go beyond the information in the data tables, SiriusBI employs a knowledge retrieval function to extract relevant information (e.g., actions taken in similar scenarios) from the knowledge management module introduced in Section 4.1, thereby assisting in addressing the current task.
Insight Summary. Insight summary is a comprehensive tool to meet various insight requirements. It automatically generates data overviews, trend analyses, anomaly diagnostics, and summary recommendations based on the characteristics of the given data table as well as the outputs from previous insight tools.
Compared with traditional BI systems and the most advanced ChatBI systems, the data insight module proposed in SiriusBI offers users enhanced data analysis capabilities. This module supports a wide range of practical application scenarios, enabling users to achieve more usable and reliable data analysis results. In addition, it addresses complex and challenging data analysis tasks, meeting the growing demand for dependable analytical solutions.
@@ 5. Experiments and Results
To evaluate our framework, in this section, we conduct an in-depth investigation from two insights: 1) On the general NL2SQL task, SiriusBI outperforms the state-of-the-art baselines on both SRD and MRD datasets; 2) The end-to-end design of the SiriusBI system is rational, where the individual performance of each module demonstrates significant superiority.
@@ 5.1. Experiment Setups
Datasets. To demonstrate the overall performance of SiriusBI and the effectiveness of its various modules, we conduct comprehensive experiments using datasets from multiple tasks. Detailed descriptions of the datasets used for each task are provided below:
NL2SQL. Although the Spider (Yu et al., 2018) and Bird (Li et al., 2023) datasets are widely used for NL2SQL tasks, their default SRD settings and the gap with industrial scenarios make it difficult to accurately evaluate NL2SQL tasks in actual BI scenarios. Therefore, we follow the previous approach (Lian et al., 2024) and select high-quality and reliable data from real application scenarios to construct a muilt-round dialogue (MRD) dataset. The dataset contains 55 sets of queries, where each set consists of 2 to 5 rounds. In total, the dataset contains 160 queries spanning 35 columns, accompanied by SQL written by human experts for reference. Among these queries, 10 sets do not fully include all metrics and dimensions, while another 10 sets exhibit confusion in metrics or dimensions. While most baselines are proposed for single-round dialogue (SRD) scenario, to ensure fair comparison, we also supplement the missing information in the MRD dataset and add 22 more single-round queries, thereby creating a SRD dataset with totally 182 queries for evaluation purposes.
Knowledge Management. To evaluate the knowledge management module, we focus on two main aspects: 1) Assessing its impact on NL2SQL tasks; 2) Evaluating the performance of knowledge retrieval. In the latter evaluation, we utilize the knowledge base corresponding to the MRD dataset, which contains knowledge across six dimensions: table, column, value, term, udf, and alias, totaling 7704 pieces of knowledge.
@@
Evaluation Metrics. In NL2SQL tasks, we follow previous work (Lian et al., 2024) and applies the Useful Execution Accuracy (UEX) metric. While other metrics like Execution Accuracy and Exact Match require the generated query matches the ground-truth query precisely, UEX recognizes a query as correct if the execution results of the generated query align with the actual results.
For tasks beyond NL2SQL within SiriusBI, such as knowledge retrieval in knowledge management, comprehensive analysis in data insight, we employ the retrieval recall rate and the execution accuracy rate, respectively.
Baselines. We compare the NL2SQL capability of SiriusBI with three state-of-the-art baselines as below:
DIN-SQL (Pourreza and Rafiei, 2023) breaks down the NL2SQL task into discrete sub-tasks, providing specific prompts for each sub-task to GPT-4o through the completion of each component.
MAC-SQL (Wang et al., 2023) employs a multi-agent framework to handle NL2SQL tasks, involving three main agents: selector, decomposer, and refiner. These agents work together to enhance the accuracy and efficiency of SQL generation.
MRD-SQL is a reproduction of previous work (Lian et al., 2024). We reproduce the basic MRD system integrated with an SQL generation model. To ensure fair comparison, we use the same base LLM as in SiriusBI.
SiriusBI employs all modules introduced in this paper. For evaluation of NL2SQL, the data insight module is excluded. SiriusBI processes user inputs through dialogue analysis and intention querying before generating SQL queries. We use “one-step” and “two-step” to distinguish between the different SQL generation methods in SiriusBI.
LLM Settings. In SiriusBI and MRD-SQL, we utilize the Hunyuan-70B-SQL, specifically designed for SQL scenarios, developed by Tencent’s Hunyuan LLM group. For other baselines, we employ the most advanced OpenAI API (Ope, [n.d.]) as the foundational LLM, specifically the ”gpt-4o-2024-08-06” version, to ensure optimal performance in all compared baselines.
Training Setting. For LLM fine-tuning in one-step SQL generation method, we use four nodes equipped with a total of 32 NVIDIA-A100-80GB GPUs, setting the training to 3 epochs, a learning rate of 3e-5, a batch size of 2, and a sequence length of 8192 tokens. The fine-tuning is conducted on a private dataset consisting of over 9,000 entries collected from industrial scenarios.
5.2. Experiment with NL2SQL
@@
In this experiment, we primarily evaluate the performance of various baselines on the SRD and MRD datasets. Note that, as an analysis module following NL2SQL, the data insights module do not participate in this part.
5.2.1. Results on SRD Dataset
We first conduct experiments on SRD dataset and present the results in Table 2. Our findings indicate that SiriusBI with both two SQL generation methods show a clear improvement over other baselines on the dataset. Specifically, when compared to competitive baselines such as GPT-4o and its derivatives (DIN-SQL and MAC-SQL), SiriusBI achieves an improvement of at least 40% in UEX. In comparison to the SOTA ChatBI implementation, MRD-SQL, the two-step and one-step variants of SiriusBI achieve improvements of 17.34% and 21.39%, respectively, which highlights the effectiveness of SiriusBI NL2SQL pipeline.
Interestingly, there is a relatively slight performance difference between two variants of SiriusBI. Concretely, the one-step method, which has been carefully fine-tuned on domain-specific data, gets 4.04% higher in UEX than the SFT-free two-step method. However, it is worth noting that the data for SFT took two weeks of annotation by four experts, and SFT itself also took 12 hours. Therefore, we recommend that users seeking higher accuracy use the one-step approach, while those prioritizing faster deployment efficiency should opt for the two-step approach.
| Baseline | UEX |
| DIN-SQL | 30.06% |
| MAC-SQL | 36.42% |
| MRD-SQL | 61.85% |
| SiriusBI two-step | 79.19% |
| SiriusBI one-step | 83.24% |
5.2.2. Results on MRD Dataset
@@
Next, we evaluate SiriusBI on the MRD dataset and present the specific experimental results in Table 4. Our findings reveal that single-round methods perform poorly on the MRD dataset, where the UEX falls below 16%. In contrast, the two-step and one-step variants of SiriusBI achieve UEX scores of 55.00% and 62.50%, respectively, demonstrating the superiority of SiriusBI in real-world NL2SQL tasks when compared to the state-of-the-art multi-round design, MRD-SQL.
Remarkably, we also investigate the impact of SFT on the MRD dataset. By comparing the performance of SiriusBI one-step and MRD-SQL with and without fine-tuning, we observe that SFT yields significant performance gains at the cost of higher resource investment. Furthermore, we find that though SiriusBI two-step does not incorporate SFT, it still improves SiriusBI one-step without SFT by 5.62% through effective retrieval of domain knowledge. This leads to a similar conclusion that users can choose the one-step method for higher accuracy and the two-step method for faster deployment.
5.2.3. In-production Results.
Besides experiments on datasets collected from industrial scenarios, we also evaluate SiriusBI in production. So far, SiriusBI has been deployed in Tencent’s finance, advertising and cloud sectors. For Tencent Finance and Advertisement, we apply the one-step method by fine-tuning LLMs on elaborately collected domain-specific datasets, achieving in-production accuracies are 97% and 89%, respectively. In the case of Tencent Cloud where SiriusBI is available to business clients, we apply the two-step method to reduce deployment costs for users. The average accuracy across 14 clients are 91%. Interestingly, we find that the in-production results are higher than those reported on the SRD and MRD datasets. We attribute this to the differences in problem distribution. In industrial scenarios, users’ questions may be simpler, while the datasets include high-quality and complex questions, which are more suitable for assessing the system’s specialized capabilities.
| Baseline | UEX |
| DIN-SQL | 11.25% |
| MAC-SQL | 15.63% |
| MRD-SQL w/o SFT | 22.50% |
| MRD-SQL | 28.13% |
| SiriusBI two-step | 55.00% |
| SiriusBI one-step w/o SFT | 49.38% |
| SiriusBI one-step | 62.50% |
5.3. Systematic Deep Analysis Experiments
In addition to NL2SQL, we also conduct a deeper analysis of the effectiveness of various components in SiriusBI. In this section, we primarily focus on analyzing modules related with MRD-Q, knowledge management and data insights.

### v2

@@ 3.4.3. One-Step SQL Generation
To improve the performance of supervised fine-tuning and reduce annotation efforts, we have developed an automated and iterative data preparation workflow guided by manually labeled data and identified bad cases. The overall pipeline is illustrated in Figure 7, consisting of two main modules: data generation and data augmentation.
Data Generation. To perform data generation, we need to utilize the SQL logs from the relevant scenarios and the corresponding schema information of the databases. SQL logs are the actual SQL code executed by users, representing the most authentic application data. However, since these are directly coded by users in the compiler, we need to use data reverse engineering to capture the users’ true intentions. The data reverse engineering functionality is primarily implemented through LLM-based prompts. The input information includes SQL, the corresponding schema information, and proprietary domain knowledge base information. On the other hand, we generate some question sets using the schema combined with LLM, and then construct some query-SQL pairs through a general SQL generation model. The query-SQL pairs obtained through these two methods will enter the data quality evaluation module. This module mainly uses the locally deployed Qwen2.5-72B (Yang et al., 2024) model for filtering, combined with manual evaluation to select the final high-quality training set. The manually annotated data will serve as the In-Context Learning (Dong et al., 2022) (ICL) demonstrations used in the aforementioned process. The specific data generation pipeline can be seen in Figure 7(a).
Data augmentation. This stage enhances data quality through two key approaches: diversity expansion and error-based negative example injection. First, we use Self-Instruct (Wang et al., 2023a) with prompts to generate a broader variety of SQL queries from seed data, significantly increasing the dataset’s size and diversity. These queries are then evaluated by a LLM (Zhang et al., 2023b), and only high-quality examples are selected for training. Second, we introduce negative examples by randomly rewriting cases from the enhanced dataset, based on error categories provided by the client. From experience, negative examples typically account for 5% of the training set, helping improve the model’s robustness by teaching it to recognize and avoid common mistakes. Together, these strategies ensure a diverse, challenging dataset that enhances the model’s accuracy and reliability.
Once the data generation and augmentation phases are complete, the refined dataset is used to train the model. If necessary, the pipeline can be iterated to further refine the dataset based on additional error analysis or user feedback, ensuring continuous improvement in data quality. This iterative refinement continues until the dataset meets the desired standards or until the project reaches the budget constraints specified by our clients.
Comparison with the previous work on data augmentation for NL2SQL systems. Early efforts in this field (Wu et al., 2021; Weir et al., 2020) were generally simple, often relying on basic templates covering only a single table, heuristic methods such as synonym replacement, or traditional deep learning models like RNNs to construct NL-SQL pairs. These approaches greatly limited the quality of the generated data. More recent work, such as ScienceBenchmark (Zhang et al., 2023a), is closer to our method, proposing an automatic data augmentation system to expand training data. The main differences are: (1) Seed Data Source: We use real SQL logs combined with data reverse engineering and NL-to-SQL translation to obtain Query-SQL pairs, while ScienceBenchmark relies on manually created data. (2) Augmentation Goals: Besides increasing data diversity, we focus on improving model robustness by injecting noise and negative examples; ScienceBenchmark mainly aims to enlarge dataset size to formalize a high-quality benchmark. (3) Augmentation Approach: We directly prompt LLMs to rewrite Query-SQL pairs, whereas ScienceBenchmark uses a multi-step pipeline of template extraction, population, NL question generation, and candidate selection.
@@ 3.4.4. Two-Step SQL Generation
To address the need for cross-domain generalization in real-world scenarios and the frequent occurrence of cold-start problems in business applications, we propose a novel two-stage SQL generation paradigm. This paradigm divides the SQL generation process into two stages: Semantic Intermediate Representation (SIR) and SQL Generation.
In the first stage, the construction of SIR focuses on abstracting user intent and data semantics, transforming complex natural language queries into a standardized intermediate form that is easy to handle. The rewriting process primarily relies on in-context learning (Dong et al., 2022) to dynamically reconstruct user inputs. Additionally, we introduce a Chain-of-Thought (CoT) (Wei et al., 2022) strategy to explicitly guide the model in step-by-step reasoning about the latent intent and constraints of the query, thereby enhancing the quality of the rewritten results.
Take the query “Analyze consumption of Tencent Cloud BI in 2023” as an example. We first extract domain-specific knowledge which includes explanations of terms (“Cloud BI” in this case) and top- demonstrations exhibiting highest similarities with the embeddings of the query as few-shot (Wang et al., 2020b) references. The input information in the SIR phase includes field descriptions retrieved from the knowledge base as well as few-shot demonstrations. The output of the SIR phase is shown below:
In the second stage, we use a code-rich LLM for SQL generation. The ”Rewriting Query” and the schema information corresponding to the user query together form the prompt. By first constructing and refining user queries before generating SQL, this two-step approach ensures high accuracy while enabling seamless migration across diverse industrial scenarios.
@@ 4. Experiments and Results
To evaluate our framework, in this section, we conduct an in-depth investigation from two insights: 1) On general NL2SQL tasks, SiriusBI outperforms the state-of-the-art baselines on both SRD and MRD datasets; 2) The end-to-end design of SiriusBI is rational, where the individual performance of each module demonstrates significant superiority.
@@ 4.1. Experiment Setups
4.1.1. Datasets
To demonstrate the overall performance of SiriusBI and the effectiveness of its various modules, we conduct comprehensive experiments using both academic benchmarks and industrial-level datasets on multiple tasks. We have released our carefully curated academic benchmark, MRD-BIRD. Due to data privacy concerns, we provide partially anonymized versions of two industrial-level datasets. They are publicly available at https://github.com/TencentBigData-SiriusAI/SiriusBI. Detailed descriptions of the datasets used for each task are provided below:
SRD NL2SQL. To validate the effectiveness of the SQL generation module, we utilize the BIRD (Li et al., 2023), a challenging large-scale database NL2SQL evaluation benchmark aimed at narrowing the gap between academic research and practical applications. It includes 95 large databases and high-quality (user question, SQL) pairs covering 37 professional fields. Besides, we collect 182 SRD queries spanning 10669 tables and involving 43 columns from 5 different business scenarios, thereby creating a SRD dataset termed as SRD-Industry.
MRD NL2SQL. This task better simulates real-world BI scenarios by evaluating both intent understanding and SQL generation in multi-round dialogue (MRD) settings. Due to the lack of a standard MRD NL2SQL benchmark, we expand part of the SRD dialogues from the BIRD dataset into 3–5 round dialogues. Specifically, we use an LLM to decompose complex SQL into turn-level queries, then generate corresponding user questions that reflect natural MRD conversation style—building on prior context without repetition. We also create MRD-Industry, a new dataset matching the database scale of SRD-Industry, with 55 query sets spanning 2–5 rounds each (160 queries over 35 columns). To increase complexity, we select 10 sets and remove certain metrics and dimensions, and select another 10 sets in which we inject confusion by replacing metrics or dimensions with ambiguous alternatives.
Knowledge Management. To evaluate the knowledge management module, we focus on two main aspects: 1) Assessing its impact on NL2SQL tasks; 2) Evaluating the performance of knowledge retrieval. In the latter evaluation, we utilize the knowledge base corresponding to the MRD-Industry dataset, which contains knowledge across six dimensions: table, column, value, term, udf, and alias, totaling 7704 pieces of knowledge.
@@ 4.1.1. Datasets
4.1.2. Evaluation Metrics
For NL2SQL tasks on the BIRD dataset, we adhere to the standard evaluation metrics as outlined in (Li et al., 2023), utilizing Execution Accuracy (EX) and Valid Efficiency Score (VES). For NL2SQL tasks on our internal datasets, we follow previous works (Lian et al., 2024; Floratou et al., 2024) and employ the Useful Execution Accuracy (UEX) metric, which considers a query correct if its intent aligns with that of the golden query. This makes UEX more suitable for business intelligence scenarios.
4.1.3. Baselines and Model Configurations
We compare the NL2SQL capability of SiriusBI with three state-of-the-art baselines as below: (1) DIN-SQL (Pourreza and Rafiei, 2023); (2) MAC-SQL (Wang et al., 2023b); (3) MRD-SQL (Lian et al., 2024); (4) SiriusBI. For the baseline models, we followed their publicly available usage instructions as closely as possible, including base models and prompt designs. For DIN-SQL and MAC-SQL, we used their original configurations and prompts without modification. We reproduce the NL2BI module of the previous work (Lian et al., 2024) as MRD-SQL. In our method and MRD-SQL, we use Qwen2.5-Coder-32B (Hui et al., 2024) as the base model for two-step SQL generation on SRD and MRD datasets. For one-step SQL generation on the BIRD dataset (and MRD-BIRD), we use XiYanSQL-QwenCoder-32B (Gao et al., 2024). This choice is motivated by its demonstrated state-of-the-art performance on this benchmark, which obviates the need for additional fine-tuning. For industrial datasets, we fine-tuned Qwen2.5-Coder-32B to better fit domain data. Specifically, we use four nodes equipped with a total of 32 NVIDIA A100-40GB GPUs, setting the training to 3 epochs, a learning rate of 3e-5, a batch size of 2, and a sequence length of 8192 tokens. The fine-tuning dataset comprises over 9,000 entries collected from real-world industrial scenarios. Overall, we ensured fair comparison by aligning baseline setups with their original methods and optimizing prompts within their intended use.
4.2. End-to-End NL2SQL Performance
@@
In this experiment, we evaluate the end-to-end NL2SQL performance of various baselines on the SRD and MRD datasets. Note that, as an analysis module following NL2SQL, the data insight module is not included in this part.
4.2.1. Results on SRD Datasets
We conduct experiments on two SRD datasets—BIRD’s dev set and SRD-Industry—with results summarized in Table 1. Our method, SiriusBI, employing both one-step and two-step SQL generation, consistently outperforms baseline approaches. The improvement is especially pronounced on the SRD-Industry dataset, where SiriusBI achieves at least a 40% increase in UEX. This significant gain is attributed to SiriusBI ’s explicit design to tackle real-world challenges such as ambiguous fields and wide tables, which general-purpose models often struggle with. We exclude MRD-SQL from this comparison since it differs from SiriusBI only in the multi-round dialogue analysis module, leading to identical performance on SRD datasets. Additionally, the one-step variant of SiriusBI, fine-tuned on domain-specific data, surpasses the two-step method by 4.05% in UEX. This suggests that supervised fine-tuning provides a measurable advantage in handling domain-specific queries, further enhancing its performance.
Methods | BIRD | SRD-Industry | |
EX(%) | VES(%) | UEX(%) | |
DIN-SQL + GPT-4o | 50.72 | 58.79 | 30.06 |
MAC-SQL + GPT-4o | 57.56 | 58.04 | 36.42 |
SiriusBI two-step | 65.32 | 67.87 | 79.19 |
SiriusBI one-step | 68.97 | 70.89 | 83.24 |
4.2.2. Results on MRD Datasets
@@
Next, we evaluate SiriusBI on two MRD datasets: MRD-BIRD and MRD-Industry, and present the experimental results in Table 2. Our findings reveal that NL2SQL methods designed for single-round dialogue perform poorly on MRD datasets. Specifically, the UEX scores of DIN-SQL and MAC-SQL on the MRD-Industry dataset fall below 16%, highlighting their limitations in handling multi-round dialogue scenarios. In contrast, the two-step and one-step variants of SiriusBI achieve UEX scores of 55.00% and 62.50%, respectively, demonstrating the consistent superiority compared to MRD-SQL in real-world NL2SQL tasks.
Methods | MRD-BIRD EX(%) | MRD-Industry UEX(%) |
DIN-SQL + GPT-4o | 41.69 | 11.25 |
MAC-SQL + GPT-4o | 42.35 | 15.63 |
MRD-SQL two-step | 46.98 | 22.50 |
MRD-SQL one-step | 50.61 | 28.13 |
SiriusBI two-step | 48.03 | 55.00 |
SiriusBI one-step | 51.14 | 62.50 |
It is worth noting that on the MRD-BIRD dataset, the performance gap between SiriusBI and MRD-SQL is narrower. Specifically, the one-step variant of MRD-SQL achieves a higher EX score than the two-step variant of SiriusBI. We attribute this phenomenon to the absence of user interaction for intent clarification in the experiments on MRD-BIRD, which diminishes the advantage of SiriusBI. However, by incorporating domain-specific knowledge during intent clarification, SiriusBI still outperforms MRD-SQL when utilizing the same SQL generation method, underscoring the effectiveness of our approach in leveraging domain-specific insights to enhance performance.
4.2.3. In-production Results
In addition to experiments on industrial datasets, we evaluate SiriusBI in production across multiple Tencent business domains. Before deployment, we apply the proposed data-condition-driven dynamic strategy switching mechanism to select the most suitable SQL generation approach based on each domain’s data characteristics. For Tencent Finance and Advertising, clients provide sufficient labeled data () and the source and target domains are identical, resulting in a semantic similarity . Under these conditions, the mechanism selects the one-step fine-tuning strategy, achieving in-production accuracies of 97% and 93%, respectively. Conversely, Tencent Cloud serves diverse users with limited labeled data and low domain similarity between our available labeled data. Thus, the mechanism chooses the two-step strategy, which attains an average accuracy of 96% across 14 randomly selected clients.
4.2.4. Discussion
We observe that the in-production results are higher than those reported on the SRD and MRD datasets. This is mainly due to three factors: 1) Knowledge Completeness: Although (MRD-)BIRD datasets provide extra knowledge, their coverage is limited compared to our comprehensive internal knowledge base, which better supports accurate SQL generation in real scenarios. 2) Data Distribution and Query Complexity: Real-world user queries in specific applications tend to be more uniform and simpler, allowing fine-tuned models to achieve higher accuracy. In contrast, SRD and MRD contain diverse and complex questions from multiple domains, making them more challenging benchmarks. 3) Iterative Feedback and Model Refinement: In-production models benefit from multiple rounds of feedback and continuous improvement guided by real failure cases (see Section 3.4.3). SRD and MRD results, however, reflect single-cycle training without such iterative updates, limiting their performance. In summary, richer knowledge, simpler real-world queries, and ongoing model refinement explain the higher accuracy observed in production compared to benchmark datasets.
4.3. Detailed Analysis